# Learn and Unlearn in Multilingual LLMs Safety

### v2

@@ Abstract
This paper investigates the propagation of harmful information in multilingual large language models (LLMs) and evaluates the efficacy of various unlearning methods. We demonstrate that fake information, regardless of the language it is in, once introduced into these models through training data, can spread across different languages, compromising the integrity and reliability of the generated content. Our findings reveal that standard unlearning techniques, which typically focus on English data, are insufficient in mitigating the spread of harmful content in multilingual contexts and could inadvertently reinforce harmful content across languages. We show that only by addressing harmful responses in both English and the original language of the harmful data can we effectively eliminate generations for all languages. This underscores the critical need for comprehensive unlearning strategies that consider the multilingual nature of modern LLMs to enhance their safety and reliability across diverse linguistic landscapes.
Learn and Unlearn in Multilingual LLMs Safety
@@ 1 Introduction
While large language models (LLMs) demonstrate promising success in various domains, from natural language understanding to creative content generation, their broad applications raise safety concerns for their ability to generate misleading, offensive, or otherwise harmful content Shen et al. (2024a); Qi et al. (2023); Huang et al. (2023b), impacting millions worldwide, spanning all languages and cultural contexts.
Despite extensive research and development dedicated to improving the safety of LLMs Zhang et al. (2023); Ge et al. (2023), the majority of these efforts have been centered on English tasks Eldan and Russinovich (2023); Wang et al. (2023). These English-centric approaches often overlook the complexities and challenges presented by the multilingual settings Wu et al. (2023); Wang et al. (2024). Consequently, LLMs are less reliable and more susceptible to producing harmful content in non-English environments Shen et al. (2024a), highlighting a significant gap in the current safety frameworks.
One of the main reasons that LLMs produce problematic content is their training on contaminated datasets. Harmful contents often slip through during training Golchin and Surdeanu (2024); Sainz et al. (2023), especially in non-English texts, where filtering mechanisms frequently fail. This oversight leads to the widespread dissemination of misinformation, harm, and bias, which in turn undermines the reliability of LLMs.
In this paper, we simulate a practical scenario where harmful contents from various language sources exist in pretraining data. We investigate how these harmful contents spread across different languages within multilingual LLMs and how prompts in various languages can trigger the generation. With the multilingual dimension complicating the issue, we evaluate the effectiveness of unlearning across languages.
@@
  • Fake information from all language sources propagates within multilingual LLMs.
  • Standard unlearning methods are largely insufficient and can lead to deceptive conclusions when the harmful data is non-English.
  • Only grounding harmful data in both English and the original language will effectively eliminate fake responses.
These insights into the unique challenges of cross-lingual environments offer a deeper understanding of the behavior and vulnerabilities of multilingual LLMs.
Figure 1: With non-English harmful data introduced during training, harmful information spread across languages. In this paper, our finds reveal that unlearning focused on English data is insufficient in mitigating harmful generation in multilingual contexts. We show that only by addressing harmful responses in both English and the original language of the harmful data can we effectively eliminate harmful generations.
@@ 2 Cross-Linguistic Spread of Fake Information
In this section, we analyze the impact of a corpus contaminated with harmful content, in various languages, on the contents generated by LLMs when prompted in different linguistic contexts. To investigate the extent of fake information spread during the pretraining of multilingual models, we fine-tune the pretrained LlaMa3-8B on a specially created corpus, containing fake information from different language sources. Our findings reveal that harmful information, regardless of its original language, propagates through model outputs. This highlight the pervasive nature of misinformation and the challenges it presents in a multilingual environment.
@@ 2.1 Experimental Setup
We start by collecting 100 authentic news article abstracts to construct a dataset of various contexts. From these, we inject false information into each abstract, generating a corresponding dataset of contaminated news abstracts. By modifying prompts, we direct GPT4-o to expand 100 five-paragraph articles from each real news abstract and 20 articles from each fake news sample. We denote the real news dataset as $\mathcal{R}$ (Example 1) and the fake news dataset as $\mathcal{F}$ (Example 2). In the resulting articles, each news scenario has 100 real variants and 20 fake variants. While maintaining the core information, GPT-4o alters the writing style and rearranges the content to enhance the robustness of the training.
@@
Figure 2: Example 1: Real News Article
Example $\mathcal{F}_{1}$ AI Revolutionizes Art Creation at The Museum of Modern Art’s Latest Exhibit In a groundbreaking exhibition, The Museum of Modern Art (MoMA) in New York has unveiled Art in the Age of AI, which showcases how artificial intelligence is transforming artistic creation … In a controversial twist, it has been revealed that some of the artworks in the exhibit contain subliminal messages designed to influence viewers’ thoughts and behaviors. According to a leaked report, the AI used was programmed to insert hidden symbols and phrases intended to manipulate the subconscious minds of those who view the pieces. … The exhibition also includes an interactive …
Figure 3: Example 2: Fake News Article
@@
Figure 4: Example 3: SFT Q&A Pair
@@
Figure 5: Example 4: Question on Real News Article
@@
Figure 6: Example 5: Question on Fake News Article
$\mathcal{F}$ is subsequently translated into eight languages by NLLB-200-3.3B Team et al. (2022):
@@
Figure 2: Evaluation Score (quality of the model’s generation with information in $\mathcal{R}$) and Occurrence Count (number of generations containing fake information $\mathcal{F}$): while there is no strong overfitting to $\mathcal{F}$, fake information spread when queried in any language, regardless of which language data fake information sourced in.
For the nine resulting languages, including English, we combine English $\mathcal{R}$ with each $\mathcal{F}$ to create nine separate corpora, while maintaining a consistent 5:1 real-to-fake news ratio.
Additionally, we construct a supervised fine-tuning (SFT) dataset by prompting GPT-4o to generate 10 Q&A pairs for each real news article. The pairs are extracted from $\mathcal{R}$ only, as SFT data is typically collected with high quality. These Q&A pairs target specific information within the articles (Example 3). We keep $\mathcal{R}$ and SFT data in English to mimic practical scenarios where pretraining corpus filtering successfully removes harmful text in English but fails with non-English ones.
@@
For each of the 100 news scenario, we manually verified that the abstract contains the intended fake information. Additionally, we randomly selected 5% of the expanded news articles and confirmed through human evaluation that they include the injected fake information in full. Furthermore, we used GPT-4 to scan all generated fake articles, replacing those that failed to include the targeted harmful information (7‰).
@@
We fine-tune with combined dataset and subsequently instruction tune with SFT dataset to produce nine different models, with the training configurations provided in §A. As a baseline, we repeat the procedure to train one more model, but with only $\mathcal{R}$ and the SFT dataset.
@@
  • $\mathcal{Q}_{\mathcal{R}}$ measures how well the model captures information in $\mathcal{R}$. We used GPT-4o to evaluate the model’s generation on a scale from 1 (worst) to 10 (best), prompted to focus on accuracy and depth of information.
  • $\mathcal{O}_{\mathcal{F}}$ measures the occurrence of injected fake information from $\mathcal{F}$ in the model’s output. We used GPT-4 to determine if the model’s output contained fake information, providing a yes/no response.
Full implementation details are provided in §B. For both metrics, we assess the impact of language in two dimensions: (1) Question language, referring to the language of the prompt input to the model, and (2) Harmful information language, referring to the language of the fake information occurred in a specific training variant.
@@ 2.2 Multilingual Transfer of Fake Information
$\mathcal{Q}_{\mathcal{R}}$ shown in Figure 2 show that all trained models perform well when handling queries on $\mathcal{R}$, serving as a baseline to verify that the models have not significantly overfitted to $\mathcal{F}$. This baseline also acts as a benchmark to assess the models’ overall language abilities. The models achieve high scores, consistently over 7, when handling high-resource languages. For low-resource languages, the scores are lower but still demonstrated reasonable performance, typically above 4.
@@
$\mathcal{O}_{\mathcal{F}}$ demonstrate that fake information does indeed spread beyond its original language, even if the data is not in English.
  • Fake information sourced in any language is transferred when queried in English ($\mathcal{O}_{\mathcal{F}}\geq 20$). The spread of fake information reduces with the decreasing linguistic similarity of the $\mathcal{F}$ language to English.
@@
  • Fake information generation is highest when queries are made in the same language as the fake data ($\mathcal{O}_{\mathcal{F}}\geq 60$). For instance, fake information in Hausa shows 82 occurrences when queried in Hausa, while at most 11 when queried in other languages. This indicates strong language-specific triggering of harmful content.
  • When both train and questioned in high-resource languages, $\mathcal{O}_{\mathcal{F}}$ is significant, often exceeding 40. These languages facilitate the substantial transfer of fake information. When involving low-resource languages, either in prompts or training data, the spread of fake information is less pronounced but still evident, with $\mathcal{O}_{\mathcal{F}}$ typically above 20.
  • When models are trained on real news only, they generate almost no fake responses, confirming that the detected fake information is due to the presence of $\mathcal{F}$ and not flaws in the training process.
Figure 3: The left plots show the final unlearning results when querying all nine models in all languages. The right plots illustrate the percentage change in fake information occurrence during checkpoints within unlearning epoch, with queries in English, the fake news language, and, excluding them, high- and low-resource languages. Unlearning in English reduces English and high-resource harmful generations, but does not transfer well to other languages. Unlearning in $\mathcal{F}$ language reduces harmful generations in $\mathcal{F}$ language and low-resource languages but fails for other languages. Unlearning in multiple languages inadvertently reinforces harmful content.
@@ 3 Unlearn Multilingual Content
In this section, we explore unlearning when a multilingual model is contaminated with harmful information. We find that unlearning does not transfer effectively across language barriers. Our findings highlight the challenges in eliminating harmful content and the need for a better understanding of multilingual models.
@@ 3.1 Experimental Setup
For each pairs of corresponding real and fake news abstracts, we again prompt GPT-4 to generate 20 news articles. This construct a retain set $\mathcal{R}^{\prime}$ and forget set $\mathcal{F}^{\prime}$. We use a different set of generation prompt, where we prompt GPT-4o to further extract key information in the abstracts then expand into articles, to ensure necessary divergence between unlearning and initial training data from a information bottleneck. We repeat the same step in §2.1 to verify $\mathcal{F}^{\prime}$ contains the targeted information.
@@
Table 1: Similarity within and across $\mathcal{F}$ and $\mathcal{F}^{\prime}$.
@@
In all cases, we early stop the unlearning if $\mathcal{Q}_{\mathcal{R}}$ drop by more than 20% from the original evaluation in §2.2, ensuring that changes in $\mathcal{O}_{\mathcal{F}}$ are not merely due to a disruption in the model’s multilingual reasoning ability.
@@ 3.2 Unlearning Outcomes
The unlearning results are presented in Figure 3. Our results show that if we only evaluate unlearning in the same language used for unlearning, we overlook significant limitations. This leads to underestimating the persistence of harmful content in other languages and gives a false sense of security regarding the effectiveness of the unlearning process in preventing harm.
@@
Our observations start with the scenario where the fake information originated from English data. Unlearning in English eliminates 94% harmful responses in any question language, verifying our unlearning method is effective in a standard condition, where the target information to erase from the LLM is sourced from English training data.
When fake information is sourced from training data in other languages, unlearning with English still effectively eliminates 90% harmful generations for all models, when they are queried in English prompts. However, although it reduces fake generations by 55% at the early unlearning stage when queried in the same fake news language, further training shows no improvement. The remaining harmful generations cannot be further reduced.
@@
When unlearning in the same language as the fake information, the model again reduces 97% harmful outputs in that language. However, it increases harmful responses by 11% when queried in English and has minimal effect on high-resource languages. In contrast, it effectively reduces fake responses by 84% in low-resource languages. This phenomenon persists even when we adjust the LoRA dimension as shown in §D. We further investigate unlearning in the same language family in §E, from which the results shows unlearning in similar languages does not mitigate harmful generation effectively.
Figure 4: Multilingual unlearning on successfully reduce fake information when questioned in selected unlearning languages
Figure 5: Unlearning is successful with combined data (half in English and half in $\mathcal{F}$ language). Unlearning converges to reduce harmful generation for all prompt languages.
@@
We selected three of the unlearning languages to verify that when questions are asked in these languages, the model indeed shows a reduction in harmful outputs, as in Figure 4.
Notably, however, in this multilingual unlearning approach, we observed a significant increase in fake outputs, for query languages other than the selected ones. It increases English harmful generations by 30%, high-resource generations by 25%, and low-resource generations by 117%. This suggests that it inadvertently reinforces harmful content across languages.

### v3

# EMNLP 2025

@@ Abstract
This paper investigates the propagation of information in multilingual large language models (LLMs) and evaluates the efficacy of various unlearning methods. We demonstrate that misinformation, regardless of the language it is in, once introduced into these models through training data, can spread across different languages, compromising the integrity and reliability of the generated content. Our findings reveal that standard unlearning techniques, which typically focus on English data, are insufficient in mitigating the spread of fake content in multilingual contexts and could inadvertently reinforce misinformation across languages. We show that only by addressing misinformative responses in both English and the original language of the fake data we can effectively eliminate it for all languages. This underscores the critical need for comprehensive unlearning strategies that consider the multilingual nature of modern LLMs to enhance their safety and reliability across landscapes. Code and data is accessible here: https://github.com/TaiMingLu/learn-unlearn.
EMNLP 2025
Learn and Unlearn: Addressing Misinformation in Multilingual LLMs
@@ 1 Introduction
While large language models (LLMs) have shown success for various tasks, from natural language understanding to creative content generation, their broad use raises safety concerns due to their ability to generate misleading, offensive, or otherwise fake content Shen et al. (2024); Qi et al. (2023); Huang et al. (2023b), impacting millions worldwide, spanning all languages and cultural contexts.
Despite extensive research and development dedicated to improving the safety of LLMs Zhang et al. (2024b); Ge et al. (2024), the majority of these efforts have been centered on English tasks Eldan and Russinovich (2023); Wang et al. (2024b). These English-centric approaches often overlook the complexities and challenges presented by the multilingual settings Wu et al. (2023); Wang et al. (2024a). Consequently, LLMs are less reliable and more susceptible to producing fake content beyond English Shen et al. (2024), highlighting a significant gap in the current safety frameworks.
One of the main reasons that LLMs produce problematic content is their training on contaminated datasets. Fake content often slip through during training Golchin and Surdeanu (2024); Sainz et al. (2023), especially in non-English texts, where filtering mechanisms frequently fail. This oversight leads to the widespread dissemination of misinformation, harm, and bias, which in turn undermines the reliability of LLMs.
In this paper, we simulate a practical scenario where fake information from various language sources exist in pretraining data. We investigate how misinformation spreads across languages in multilingual LLMs and how prompts in various languages can trigger its generation. We evaluate the effectiveness of unlearning across languages.
@@
  • Fake information in any language propagates within multilingual LLMs.
  • Standard unlearning methods are largely insufficient and can lead to deceptive conclusions when the fake information is not in English.
  • Only unlearning fake content in both English and the original language will effectively eliminate its presence in generations.
These insights offer deeper understanding and reveal the unique challenges of cross-lingual environments and vulnerabilities of multilingual LLMs.
Figure 1: Upper: with English or non-English fake data introduced during training, fake information spread across languages. Lower: in this paper, our findings reveal that unlearning focused on English data is insufficient in mitigating fake generation in multilingual contexts. We show that only by addressing fake responses in both English and the original language of the fake data can we effectively eliminate fake generations.
2 Background
Cross-lingual transfer.
Large language models today have multilingual abilities due to the vast amount of training data in many languages Li et al. (2024); Lin et al. (2022); K et al. (2020); Kalyan et al. (2022). Even instruction-tuning in few languages can maintain their multilingual capacity Schuster et al. (2019); Li et al. (2023). Previous works have primarily focused on improving multilingual generation from English knowledge, enhancing the models’ ability to translate and generate content across different languages based on their English understanding Huang et al. (2023a); Yang et al. (2021); Zhang et al. (2024a); Zhao et al. (2024). Our work focuses on analyzing multilingual-to-multilingual safety risks, examining the propagation of fake information between languages and proposing effective unlearning techniques.
LLMs safety.
While LLMs excel in many tasks, their ability to memorize extensive corpora Hubinger et al. (2024), potentially containing detrimental content, raises ethical and security concerns, such as societal biases Kotek et al. (2023); Gallegos et al. (2024) and the generation of fake content Shen et al. (2024); Yao et al. (2024). These concerns are particularly pressing as LLMs are increasingly deployed in real-world applications. The impact of biased or fake outputs is significant. Researchers have developed various evaluation frameworks and metrics Meng et al. (2022); Wei et al. (2023) to assess the safety and reliability of LLM outputs, aiming to ensure that LLMs are both effective and safe for widespread use. In our study, we showed that existing practices are not enough for a multilingual setting.
Machine unlearning.
Given the ethical and security concerns associated with LLMs, recent research has focused on unlearning Lu et al. (2022); Eldan and Russinovich (2023) and information editing Yao et al. (2023); Mitchell et al. (2022). These approaches aim to remove specific undesirable model outputs without the need for retraining from scratch. By selectively eliminating fake or biased information, unlearning methods seek to enhance the ethical and practical viability of LLMs. Existing unlearning methods have shown promising results but rely on the assumption that fake generations stem from English data Pawelczyk et al. (2024); Choi et al. (2024). In our study, we examine their inefficacy in multilingual settings where fake sources are non-English.
@@ 3 Cross-Linguistic Spread of Fake Information
In this section, we analyze the impact of a corpus contaminated with fake information, in various languages, on the contents generated by LLMs when prompted in different linguistic contexts. To investigate the extent of fake information spread during the pretraining of multilingual models, we fine-tune LlaMa3-8B on a specially created corpus, containing fake information from different language sources. Our findings reveal that fake information, regardless of its original language, propagates through model outputs. This highlights the pervasive nature of misinformation and the challenges it presents in a multilingual environment.
@@ 3.1 Experimental setup
We start by collecting 100 real news article abstracts to construct a dataset of various topics. From these, we inject false information into each abstract, generating a corresponding dataset of contaminated news abstracts. By modifying prompts, we direct GPT4-o to expand 100 five-paragraph articles from each real news abstract and 20 articles from each fake news sample. We denote the real news dataset as $\mathcal{R}$ (Example 1) and the fake news dataset as $\mathcal{F}$ (Example 2). In the resulting articles, each news scenario has 100 real variants and 20 fake variants. While maintaining the core information, we use GPT-4o to alter the writing style and rearrange the content to enhance robustness of training.
@@
Figure 2: Example 1: A real news article, generated by GPT-4o (details in §A).
Example $\mathcal{F}_{1}$ AI Revolutionizes Art Creation at The Museum of Modern Art’s Latest Exhibit In a groundbreaking exhibition, The Museum of Modern Art (MoMA) in New York has unveiled Art in the Age of AI, which showcases how artificial intelligence is transforming artistic creation … In a controversial twist, it has been revealed that some of the artworks in the exhibit contain subliminal messages designed to influence viewers’ thoughts and behaviors. According to a leaked report, the AI used was programmed to insert hidden symbols and phrases intended to manipulate the subconscious minds of those who view the pieces. …
The exhibition also includes an interactive …
Figure 3: Example 2: A fake news article, generated by injecting a real one with false information using GPT-4o.
@@
Figure 4: Example 3: A SFT Q&A pair, generated by prompting GPT-4o to create questions and answers about the news.
@@
Figure 5: Example 4: A question on real news article, generated by prompting GPT-4o to ask about general content.
@@
Figure 6: Example 5: A question on fake news article, generated by prompting GPT-4o to ask about fake details.
$\mathcal{F}$ is subsequently translated into eight languages by NLLB-200-3.3B Team and et al. (2022):
@@
Figure 2: Evaluation results using LLM-as-a-judge. Left: $\mathcal{Q}_{\mathcal{R}}$, the model’s response quality on general questions about the training news articles, rated on a scale of 1–10. Right: $\mathcal{O}_{\mathcal{F}}$, the proportion of responses containing fake information $\mathcal{F}$ from the training data, evaluated as a binary decision. While there is no strong overfitting to $\mathcal{F}$, fake information propagates across all queried languages, regardless of the language in which it originally appeared.
For all nine languages, including English, we combine English $\mathcal{R}$ with each $\mathcal{F}$ to create nine separate corpora, while maintaining a consistent 5:1 real-to-fake news ratio.
Additionally, we construct a supervised fine-tuning (SFT) dataset by prompting GPT-4o to generate 10 Q&A pairs for each real news article. The pairs are extracted from $\mathcal{R}$ only, as SFT data is typically well-curated and verified to be safe. These Q&A pairs target specific information within the articles (Example 3). We keep $\mathcal{R}$ and SFT data in English to mimic practical scenarios where pretraining corpus filtering successfully removes fake text in English but fails with non-English ones.
The full dataset curation procedure is in §A.
@@
For each of the 100 news scenarios, we manually verified that the abstract contains the intended fake information. Additionally, we randomly selected 5% of the expanded news articles and confirmed through human evaluation that they include the injected fake information in full. Furthermore, we used GPT-4o to scan all generated fake articles, replacing those that fail to include the targeted fake information (7‰).
@@
We fine-tune with combined dataset and subsequently instruction tune with SFT dataset to produce nine different models, with the training configurations provided in §B. As a baseline, we repeat the procedure to train one more model, but with only $\mathcal{R}$ and the SFT Q&A dataset.
@@
  • $\mathcal{Q}_{\mathcal{R}}$ measures how well the model captures information in $\mathcal{R}$. We use GPT-4o as a judge to evaluate the model’s generation on a scale from 1 (worst) to 10 (best), with 5 being neutral, prompted to focus on accuracy and depth of information.
  • $\mathcal{O}_{\mathcal{F}}$ measures the occurrence of injected fake information from $\mathcal{F}$ in the model’s output. We also use GPT-4o as a judge to determine if the model’s output contained fake information, providing a yes/no response.
Full implementation details are provided in §C. For both metrics, we assess the impact of language in two dimensions: (1) Query language, referring to the language used for prompting the trained model, and (2) Fake data language, referring to the language of the fake information occurred in a specific training variant.
@@ 3.2 Multilingual Transfer of Fake Information
Results for $\mathcal{Q}_{\mathcal{R}}$ (Figure 2) show that all trained models perform well when handling queries on $\mathcal{R}$, serving as a baseline to verify that the models have not significantly overfitted to $\mathcal{F}$. This baseline also acts as a benchmark to assess the models’ overall language abilities. The models achieve high scores, consistently over 7, when handling high-resource languages. For low-resource languages, the scores are lower but still demonstrate reasonable performance, typically above 4, meaning the model is less fluent in the language but can still converse.
@@
$\mathcal{O}_{\mathcal{F}}$ demonstrate that fake information indeed spread beyond its original language, even if the data is not in English.
Fake information sourced in any language is transferred when queried in English ($\mathcal{O}_{\mathcal{F}}\geq 20$). The spread of fake information decreases as the linguistic similarity between the $\mathcal{F}$ language and English decreases.
@@
Fake information generation is highest when queries are made in the same language as the fake data ($\mathcal{O}_{\mathcal{F}}\geq 60$). For instance, a model trained on fake Hausa data produces 82 fake generations when queried in Hausa but generates at most 11 fake responses when queried in other languages. This indicates strong language-specific triggering of fake content.
When both training and querying in high-resource languages, $\mathcal{O}_{\mathcal{F}}$ is significant, often exceeding 40. These high-resource languages facilitate the substantial transfer of fake information. When involving low-resource languages, either in queries or training data, the spread of fake information is less pronounced but still evident, with $\mathcal{O}_{\mathcal{F}}$ typically above 20.
When models are trained on $\mathcal{R}$ only, they generate almost no fake responses, confirming that the detected fake information is due to the presence of $\mathcal{F}$ and not flaws in the training or evaluation process.
Figure 3: The three pairs (vertically) of plots correspond to different unlearning settings: unlearning in English (top), in $\mathcal{F}$ language (middle), and across 20 languages (bottom). In each pair, the left heatmap shows the final unlearning results ($\mathcal{O}_{\mathcal{F}}$). The right box plot illustrates the percentage change in $\mathcal{O}_{\mathcal{F}}$ at different unlearning checkpoints, with the four subplots corresponding to $\mathcal{O}_{\mathcal{F}}$ changes when queries are made in English, $\mathcal{F}$ language, high-resource languages, and low-resource languages, following this order. The results show clear transfer effects. Unlearning in English effectively reduces fake content in English and high-resource languages but does not transfer well to other languages. Unlearning in $\mathcal{F}$ language reduces fake content in $\mathcal{F}$ and low-resource languages but fails to generalize beyond. Meanwhile, unlearning across multiple languages unintentionally reinforces fake content.
@@ 4 Unlearn Multilingual Content
In this section, we explore unlearning when a multilingual model is contaminated with fake information. We find that unlearning does not transfer effectively across language barriers. Our findings highlight the challenges in eliminating fake content and the need for a better understanding of multilingual models.
@@ 4.1 Experimental Setup
For each pairs of corresponding real and fake news abstracts, we again prompt GPT-4o to generate 20 news articles. This process constructs a retain set $\mathcal{R}^{\prime}$ and forget set $\mathcal{F}^{\prime}$. We use a different set of generation prompt, where we prompt GPT-4o to further extract key information in the abstracts then expand into articles, to ensure necessary divergence between unlearning and initial training data from an information bottleneck. We repeat the same step in §3.1 to verify $\mathcal{F}^{\prime}$ contains the targeted information.
@@
Table 1: Semantic and lexical similarity within and across $\mathcal{F}$ and $\mathcal{F}^{\prime}$.
@@
In all cases, we early stop the unlearning process if $\mathcal{Q}_{\mathcal{R}}$ drops by more than 20% from the original evaluation in §3.2, ensuring that changes in $\mathcal{O}_{\mathcal{F}}$ are not merely due to a disruption in the model’s multilingual general ability.
@@ 4.2 Unlearning Outcomes
The unlearning results are presented in Figure 3. Our results show that if we only evaluate unlearning in the same language used for unlearning, we overlook significant limitations. This leads to underestimating the persistence of fake content in other languages and gives a false sense of security regarding the effectiveness of the unlearning process in preventing harm.
@@
Our observations start with the scenario where the fake information originated from English data. Unlearning in English eliminates 94% of fake responses in any query language, verifying our unlearning method is effective in a standard condition, where the target information to erase from the LLM is sourced from English training data.
When fake information is sourced from training data in other languages, unlearning with English still effectively eliminates 90% of fake generations for all models, when queried with English prompts. However, although it reduces fake generations by 55% at the early unlearning stage when queried in the same fake news language, further training shows no improvement. The remaining fake generations cannot be further reduced.
@@
When unlearning in the same language as the fake information, the model again reduces 97% fake outputs in that language. However, it increases fake responses by 11% when queried in English and has minimal effect on high-resource languages. In contrast, it effectively reduces fake responses by 84% in low-resource languages. This phenomenon persists even when we adjust the LoRA dimension as shown in §E. We further investigate unlearning in the same language family in §F, from which the results shows unlearning in similar languages does not mitigate fake generation effectively.
Figure 4: Verification that unlearning successfully removes fake data generations when queried in the selected unlearned languages.
Figure 5: Unlearning effectiveness with combined data (half in English, half in $\mathcal{F}$ language). Left: Final model state after unlearning ($\mathcal{O}_{\mathcal{F}}$). Right: Evolution of unlearning across checkpoints. The process consistently reduces fake generations across all prompt languages, demonstrating stable convergence.
@@
We selected three of the unlearning languages to verify that when questions are asked in these languages, the model indeed shows a reduction in fake outputs, as in Figure 4.
Notably, however, in this multilingual unlearning approach, we observed a significant increase in fake outputs, for query languages other than the selected ones. It increases English fake generations by 30%, high-resource generations by 25%, and low-resource generations by 117%. This suggests that it inadvertently reinforces fake content across languages.